OpenStudy (anonymous):

What is x 8x = 4x – 32

pooja195 (pooja195):

Subtract 4x from both sides 8-4=?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

yea what does 8-4=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Good :) now you have 4x=-32 any ideas on what to do next?

OpenStudy (anonymous):

subtract it

pooja195 (pooja195):

Not quite we need to divide it so that we can isolate x 4/4=1 right that leaves us with x x=? -32/4=?

pooja195 (pooja195):

need help?